OpenStudy (anonymous):

x^2-16 / x+4 .....what is this in simplified form ??????? x^2-16 / x+4 .....what is this in simplified form ??????? @Mathematics

OpenStudy (anonymous):

x-4

OpenStudy (anonymous):

with x#-4

OpenStudy (anonymous):

\[(x-4) \] but why ?

OpenStudy (anonymous):

\[\frac{x^2 - 4^2}{x+4} = \frac{(x-4)(x+4)}{x+4} = (x-4)\]

OpenStudy (anonymous):

why \[ x^2 - 4^2 = (x-4)(x+4) \]

OpenStudy (anonymous):

hahahahaha fool mazak achha hai....................

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

pariksha na lo hamari parbhu........

OpenStudy (anonymous):

lolz naah bhai just fooling around :P

OpenStudy (anonymous):

ok buddy now loggin off will catch u later....bbye cya

OpenStudy (anonymous):

see you sheggy :)